OpenStudy (anonymous):

a plane traveled 3246 miles with the wind in 6 hours, and 2814 miles against the wind in the same time. Find the speed of the plane and the wind

OpenStudy (anonymous):

Divide to find the two speeds.

OpenStudy (anonymous):

Then you'll have plane + wind for one and plane - wind for the other.

OpenStudy (anonymous):

Those will be your two equations to solve for the two unknowns.

OpenStudy (anonymous):

that really didnt help.

OpenStudy (anonymous):

Have you done any similar problems before?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

What don't you understand about what I told you to do?

OpenStudy (anonymous):

you were vague

OpenStudy (anonymous):

i onyl have 4 minutes can someone pleqsae help me

OpenStudy (anonymous):

\[\frac{ 3246 }{ 6 }=v_p+v_w\] \[\frac{ 2814 }{ 6 }=v_p-v_w\]

OpenStudy (anonymous):

I'm vague because I want you to do some thinking on your own, @strcaitlyn13 . I have respect for your intelligence and don't want to insult you by telling you things you already know.

OpenStudy (anonymous):

3246/(p+w) =6 2184/(p-w) =6 p+w=3246/6 p-w=2184/6 solve for p and w p+w=541 p-w=364 add 2p=905 p=905/2 subtract 2w=177 w=177/2

OpenStudy (anonymous):

And then there's always that helpful soul who will insult your intelligence and betray the sanctity of the learning process (*cough* code of conduct *cough*) by merely giving you the solution. Oh well, not everyone wants to actually be educated.

OpenStudy (anonymous):

so is it 88.5? and 354

OpenStudy (anonymous):

452.5,88.5

OpenStudy (anonymous):

Ah, all that work, and your solutions are incorrect, @gomathi ...

OpenStudy (anonymous):

Here's a hint: The solutions are whole numbers.

OpenStudy (radar):

One of the advantages of math, is the ability to verify your answers or "checking" them. In this case, let us verify the answer 452.5 mph for the airplane speed and 88.5 mph for the wind speed. when the plane was flying with the wind, the combined speed using d=rt, 3246=r6, thus rate is 3246/r or 541 mph now for check: 452.5 + 88.5 = 541.0 that checks out.

OpenStudy (radar):

When the plane is flying against the wind, you can also verify those values.

OpenStudy (anonymous):

Uh huh, and do those check out?

OpenStudy (radar):

They do indeed.

OpenStudy (anonymous):

Really? "a plane traveled 3246 miles with the wind in 6 hours, and 2814 miles against the wind in the same time. Find the speed of the plane and the wind"

OpenStudy (radar):

As show in my post the combined rate (mph) when traveling with the wind would be 3246/6 or a rate of 541 mph. This is the result of the airplane's speed plus the wind speed. or plane + wind = 541 mph. In a similar manner calculate the combined rate when flying in opposition to the wind (against the wind) that would be 2814/6 = 469 mph. This is the airplane speed minus the wind speed. let A = airplane speed and let W = wind speed A+W=541 mph A-W=469 mph using elimination method 2A = 1 0 1 0 A= 505 New "corrected" speed if plane is now 505 mph (lol) W=541-505 =36 mph a more reasonable wind speed ! I see where the value of 2184 was used instead of 2814 resulting in an error. Good catch Cliffsedge.

OpenStudy (anonymous):

I wouldn't want to fly in hurricane force winds either.